# Task Graph offloading via Deep Reinforcement Learning in Mobile Edge Computing

### v1

@@ 1 Introduction
With the rapid advance of wireless communication technologies and the proliferation of smart mobile devices, such as smartphones, tablets, and wearable devices, these lightweight smart devices have become significant terminals for mobile users (MUs) to access the Internet. Various mobile applications are gaining widespread popularity and are increasingly complex. Moreover, they are often composed of dependent tasksgeng2018energy that form the construct of a task graph, such as gesture recognition, mobile healthcare, and augment realitynovak2018ultrasound ; goudarzi2020application . These applications often have low-latency requirements, resulting in a significant surge in demand for computing resources. Nevertheless, it is difficult for smart devices to effectually support the local execution of these computation-intensive applications due to the naturally lightweight and limited computing resourcesren2019survey . The tension between computation-intensive applications and resource-limited smart devices generates a bottleneck for obtaining a high-quality user experience.
@@
Most works have focused on optimal task scheduling strategies for task graph offloading in terms of the application completion timezhang2018resource , energy consumptionsundar2018offloading , and communication servicenaouri2021novel . Due to the NP-hardness of task scheduling optimization, most studies on task graph offloading in MEC are based mainly on heuristic or approximation methodssundar2018offloading ; zhao2020offloading ; al2020task . In real-world scenarios, ECDs in the network are mainly responsible for some specialized functionsdong2022wave , such as routing, relaying, forwarding, and so on. In this case, ECDs only generally share their idle computing resources with the applications offloaded from MUs. Considering that ECDs are still relatively resource-constrained devices compared to the remote cloud, their idle computing resources are incapable of maintaining persistent performance. The computing capabilities of ECDs may vary with time making the MEC environment dynamic. However, these aforementioned works rely heavily on expert knowledge or accurate analytical models so they fail to fully adapt to the dynamic environment, resulting in the deadline violation for the mobile applications since the actual task execution may fluctuate frequentlyliu2019online ; al2020task ; zhao2020offloading . Consequently, a few works have been attracted to research the computation offloading in MEC by utilizing the novelty technologies that can adapt to the dynamic environment.
@@
A few works investigate the task offloading in MEC by utilizing DRL to adapt to the environment variationhe2017integrated ; ning2019deep ; chen2018optimized . He et al. he2017integrated improved the performance of computation offloading for vehicular networks by dynamic orchestration of network, caching, and computing resources. Ning et al. ning2019deep considered the variation of the channel state and the computation capability and proposed a DRL scheme to optimize traffic scheduling and resource allocation in vehicular networks. Chen et al. chen2018optimized took into account the time-varying network dynamics to investigate an optimal computation offloading policy in a sliced radio access network. They proposed a DRL-based offloading decision to maximize the long-term utility performance. However, these works focus the coarse-grained task offloading rather than fine-grained task graph offloadingdinh2017offloading . Coarse-grained task offloading regards an application as an indivisible whole and designs the offloading scheme based on applications’ requirements on computing resources. This scheme does not divide the mobile application into tasks from its functional point of view and may result in unexpected computation delay by unreasonable processing compositor of tasks with dependency constraints.
Many existing works apply DRL to investigate the fine-grained offloading in MEClu2020optimization ; song2022offloading ; wang2021dependent . Lu et al. lu2020optimization focused on the scenarios of the multi-service nodes and task graph applications in heterogeneous MEC. They proposed a fine-grained offloading scheme based on DRL to achieve the reduction of execution latency, monetary cost, energy consumption, as well as network usage. Song et al. song2022offloading considered the different dynamic preferences between MEC and users to investigate the offloading decision. They proposed a multi-objective optimization strategy based on DRL for task graph applications in MEC. Wang et al.wang2021dependent proposed a DRL-based task graph offloading scheme. In their proposal, an off-policy RL empowered by a Sequence-to-Sequence neural network is utilized to make the task offloading plan, while the input of this neural network is the task graph model. Although these above-mentioned works have proposed various task graph offloading strategies, they did not consider the dynamic environment of MEC. This consequently motivates us to consider the time-varying computing capabilities of ECDs to improve the performance of the task graph offloading in MEC.
@@
In this paper, we investigate the task graph offloading in MEC when considering the computation capabilities of ECDs are finite continuous values varying with time. We first formulate the task graph offloading in MEC as an optimization problem of minimizing the average application completion. Then, we model the task graph scheduling for computation offloading as a Markov Decision Process (MDP). In this MDP, the characterization of the environment is formulated as the state space, and the task scheduling decisions are abstracted into the action space. The reward with respect to MDP is defined as the benefit for the agent. To overcome the difficulties caused by the large number of state spaces, we designed a DRL algorithm based on DQN to learn the task scheduling strategy from the interaction with the environment. Particularly, the major contributions of this paper are summarized as follows.
@@
    The task graph offloading in MEC is formulated as a problem of minimizing the average application completion time.
@@
    The processing of the task graph offloading is abstracted into a sequence of discrete time steps. The interaction between the agent and environment is represented to model a MDP.
@@
    The process of task graph scheduling is modeled as a MDP. According to the processing of the task graph offloading, the state space, the action space, and the reward are formulated and a DRL algorithm based on DQN is designed to learn the scheduling strategy.
  • 4.
    By conducting extensive simulations, we confirm that the proposed task graph offloading algorithm is superior to existing strategies, in terms of reducing average makespan and deadline violation.
The remainder of this paper is organized as follows. Section 2 presents the system model and Section 3 formulates the optimization problem of the task graph offloading in MEC. Section 4 gives the RL-based scheduling mechanism, and Section 5 describes the resource allocation based on RL. Then, the simulation evaluation and conclusions are provided in Section 6 and Section 7, respectively.
@@ 2 System Mobel
A MEC system, as shown in Fig. 1, is comprised of many heterogeneous edge computing devices (ECDs) deployed at the edge of the mobile network. Let $\mathcal{M}$ express the set of ECDs in the MEC system and $M=|\mathcal{M}|$ denote the total number of ECDs. Then, we use $m$ ($\forall m\in\{1,\ldots,M\}$) to index the $m$-th ECD. These ECDs in $\mathcal{M}$ communicate with each other via wired links zhang2017resource . We indicate the transmission rate between any two ECDs as $B_{m,m^{\prime}}$, where $m\neq m^{\prime}$ and $\forall m,m^{\prime}\in\{1,\ldots,M\}$. Based on the literature zhang2015energy , $B_{m,m^{\prime}}$ can be estimated generally. Due to physical constraints, the computing resources of ECDs are limited. We assume that each ECD has only one processing element (PE) which provides the computing service for the task execution of the application following First Come First Service manner (FCFS). In practice, the processing capabilities of PEs are finite continuous values varying with timening2019deep . Moreover, the processing capability state at the next moment is related to the previous moment. We assume that the processing capabilities of PEs are stabilized over a short period when they are within a level. Therefore, the processing capabilities of PEs can be discretized and quantized into several levels, similar to the literaturening2019deep . Then, let $\delta_{m}$ express the processing capability of ECD $m$ at a state, which can be measured in millions of instructions per second (MIPS). For the scenarios in which each ECD has multiple PEs with parallel computing capability, we can add more virtual ECDs and let the data transmission cost among them be 0. Each ECD can periodically broadcast the current status information, such as its own workload and the transmission rate with other ECDs.
@@
In the MEC system shown in Fig. 1 each ECD can provide cellular communication services to many mobile users (MUs) in a specified area. We use $\mathcal{N}$ to indicate the set of all MUs that the MEC system can serve, and $N=|\mathcal{N}|$ denotes the total number of MUs. Thus, $n$ ($\forall n\in\{1,\ldots,N\}$) can index the $n$-th MU. Each MU needs to process an application composed of multiple dependent tasks. This proposal can also be easily extended to the scenarios in which a MU can process multiple applications by adding some new MUs that can only process one application to the system. As a result, $\mathcal{N}$ also represents the set of all applications and $n$ can index the $n$-th application [^1] In this paper, we will interchangeably use three terms, i.e., MU $n$, $n$-th MU and application $n$, to represent that a user application. We can model the application as a direct acyclic graph (DAG) sundar2018offloading . Then, $G_{n}(\mathcal{V}_{n},\mathcal{E}_{n})$ indicates a DAG corresponding to application $n$. $\mathcal{V}_{n}$ denotes the set of tasks. $\mathcal{E}_{n}$ is the set of data dependencies, called directed edges, among these tasks. For any task in $\mathcal{V}_{n}$, it is an atomic and indivisible component.
@@
There is a decision controller in the MEC system, which resides in some ECDs zhang2017resource , shown in Fig. 1. It can periodically collect the real-time status information broadcasted from all ECDs. MU $n$ expects to expedite the execution of its application before the deadline, so it needs to offload application data to the MEC system. To achieve this, MU $n$ must first send the offloading request to the ECD whose cellular communication service covers MU $n$. Let $m_{n}$ indicate the ECD that covers MU $n$, as shown in Fig. 1. $B_{n}^{m}$ denotes the transmission rate between ECD $m$ and MU $n$. After ECD $m_{n}$ accepts this request and receives the application data sent by MU $n$, the decision controller will extract the tasks ready for execution from ECD $m_{n}$ and push them into a ready queue $Q^{r}$. Whereafter, the decision controller can make the assignment for the tasks in $Q^{r}$ according to a scheduling strategy. Moreover, any task can only be executed on one ECD. The result data must send back to MU $n$ after all tasks of application $n$ are completed in the MEC system. As MU $n$ is located within the cellular signal coverage area of ECD $m_{n}$, the result data needs to be transmitted to MU $n$ via ECD $m_{n}$. For any MU $n$, it has the same network topology for task execution as all ECDs in this system can execute tasks offloaded from MU $n$. Hence, we only focus on the topology with one MU for simplicity, which is shown in Fig. 2. More MUs connected to the MEC system only increase the number of applications offloaded to this system, which does not affect this study. Built upon this topology with one MU, we can assume MU $n$ as the fictitious ECD which is the 00-th edge computing device in $\mathcal{M}$. To model the application data sent by MU $n$ and the result data received by MU $n$, we add two dummy tasks to each application in $\mathcal{N}$. Moreover, the two dummy tasks must be on this fictitious ECD corresponding to MU $n$. As a result, all tasks in application $n$ can be offloaded to the MEC system.
@@
Considering the above characteristics of the computing offloading, we use the triple elements $\{r_{n},d_{n},\mathcal{G}_{n}\}$ to indicate the application that MU $n$ offloads to the MEC system. Here, $r_{n}$ denotes the time of the offloading request sent by MU $n$ and $d_{n}$ is the deadline of application $n$. In the task graph $\mathcal{G}_{n}$, $I_{n}=|\mathcal{V}_{n}|$ expresses the number of tasks in application $n$. We indicate each task in $\mathcal{V}_{n}$ as $v_{ni}$ ($i\in\{0,\ldots,I_{n}\}$). Here, $i$ indexes the $i$-th task. For ease of expression, meanwhile, we use $v_{n0}$ and $v_{nI}$ to denote the two dummy tasks, respectively. For task $v_{ni}$, it has a workload $\rho_{ni}$, which is the number of instructions needed by the PE to process $v_{ni}$. Generally, $\rho_{ni}$ can be acquired by profile analytics venkataraman2016ernest . Accordingly, $\rho_{n0}$ and $\rho_{nI}$ are set to zero since $v_{n0}$ and $v_{nI}$ are the dummy tasks added to application $n$. Due to the dependency constraints among tasks, we represent the directed edge between $v_{ni}$ and $v_{nj}$ as $\varepsilon_{nij}\in\mathcal{E}_{n}$. If there exist the output data from $v_{ni}$ to $v_{nj}$, we call $v_{ni}$ the parent of $v_{nj}$. Accordingly, $v_{nj}$ is the child of $v_{ni}$. We use $p(v_{ni})$ and $c(v_{ni})$ to define the sets of $v_{ni}$’s parents and children, respectively. For the two dummy $v_{n0}$ and $v_{nI}$, especially, $p(v_{n0})$ and $c(v_{nI})$ are empty sets. Under the scheduling strategy made by the decision controller, the output data from a completed task need to be transferred to the ECD where the child tasks are located, due to the data dependencies between tasks. For example, task $v_{ni}$ is on ECD $2_{n}$ and $v_{ni}$’s child $v_{nj}$ is on ECD $3$ in Fig. 2. After ECD $2_{n}$ completes task $v_{ni}$, ECD $2_{n}$ needs to transmit the all output data to ECD $3$ via the communication link since $v_{ni}$’s child $v_{nj}$ is on ECD $3$. Let $e_{nij}$ denote the size of data transfer from $v_{ni}$ to $v_{nj}$. Especially, if both $v_{ni}$ and its child $v_{nj}$ on the same ECD $m$, the transmit data $e_{nij}$ from $v_{ni}$ can be delivered directly to $v_{nj}$ without going through the communication link.

### v4

@@ 1 Introduction
With the rapid advance of wireless communication technologies and the proliferation of smart mobile devices, such as smartphones, tablets, and wearable devices, these lightweight smart devices have become significant terminals for mobile users (MUs) to access the Internet. Various mobile applications are gaining widespread popularity and are increasingly complex JIT2217 . Moreover, they are often composed of dependent tasksgeng2018energy that form the construct of a task graph, such as gesture recognition, mobile healthcare, and augment realitynovak2018ultrasound ; goudarzi2020application . These applications often have low-latency requirements, resulting in a significant surge in demand for computing resources. Nevertheless, it is difficult for smart devices to effectually support the local execution of these computation-intensive applications due to the naturally lightweight and limited computing resourcesren2019survey . The tension between computation-intensive applications and resource-limited smart devices generates a bottleneck for obtaining a high-quality user experience.
@@
Most works have focused on optimal task scheduling strategies for task graph offloading in terms of the application completion timezhang2018resource , energy consumptionsundar2018offloading , and communication servicenaouri2021novel . Due to the NP-hardness of task scheduling optimization, most studies on task graph offloading in MEC are based mainly on heuristic or approximation methodssundar2018offloading ; zhao2020offloading ; al2020task ; meng2016approach . In real-world scenarios, ECDs in the network are mainly responsible for some specialized functionsdong2022wave ; IOT9320569 , such as routing, relaying, forwarding, and so on. In this case, ECDs only generally share their idle computing resources with the applications offloaded from MUs. Considering that ECDs are still relatively resource-constrained devices compared to the remote cloud, their idle computing resources are incapable of maintaining persistent performance. The computing capabilities of ECDs may vary with time making the MEC environment dynamic. However, these aforementioned works rely heavily on expert knowledge or accurate analytical models so they fail to fully adapt to the dynamic environment, resulting in the deadline violation for the mobile applications since the actual task execution may fluctuate frequentlyliu2019online ; al2020task ; zhao2020offloading ; wang2021dependent . Consequently, a few works have been attracted to research the computation offloading in MEC by utilizing the novelty technologies that can adapt to the dynamic environment.
@@
A few works investigate the task offloading in MEC by utilizing DRL to adapt to the environment variationhe2017integrated ; ning2019deep ; chen2018optimized ; xu2019confidence . These works consider the dynamic variation of the edge computing network to improve the performance of computation offloading. However, these works focus the coarse-grained task offloading rather than fine-grained task graph offloadingdinh2017offloading . Coarse-grained task offloading regards an application as an indivisible whole and designs the offloading scheme based on applications’ requirements on computing resources. This scheme does not divide the mobile application into tasks from its functional point of view and may result in unexpected computation delay by unreasonable processing compositor of tasks with dependency constraints.
Recently, some works apply DRL to investigate the fine-grained offloading in MECsong2022offloading ; wang2021dependent . Song et al. song2022offloading took into account the different dynamic preferences between MEC and users to solve the multi-objective optimization problem of task graph offloading in MEC. Wang wang2021dependent focused on the variation of dependent constraints between tasks to improve the Quality-of-Service (QoS) of task graphs. Although these works have proposed various task graph offloading strategies, they did not consider the variation of computing capabilities of ECDs. The dynamic variation will lead to task execution delays. Moreover, the effect will continuously accumulate until task graph applications are complete if the offloading strategy can not adapt to the dynamic variation, which leads to the deadline violation. This consequently motivates us to investigate DRL-based task graph offloading to improve user experience, considering the time-varying computing capabilities of ECDs.
@@
In this paper, we propose a task graph offloading strategy based on DRL in the MEC system. The proposed strategy can adapt to the dynamic environment when the computing capabilities of ECDs are time-varying, to satisfy application deadlines and improve user experience as much as possible. We first formulate the task graph offloading in MEC as an optimization problem of minimizing the average application completion. Based on the topological ordering of applications, we decompose the ready tasks of the arrival applications into a prioritized list. Then, we model innovatively the scheduling processing of the ready tasks as a Markov Decision Process (MDP). In this MDP, the characterization of the environment is formulated as the state space, and the scheduling decision for each ready task is abstracted into the action space. The reward with respect to MDP is defined as the benefit for the agent. To overcome the difficulties caused by the large number of state spaces, we designed a DRL algorithm based on DQN to learn the scheduling decision from the interaction with the environment. Particularly, the major contributions of this paper are summarized as follows.
@@
    The task graph offloading is decomposed into making the scheduling decision for the ready tasks of the arrival applications in a sequence of discrete time steps. The scheduling processing of the ready tasks is modeled as an MDP innovatively.
@@
    The environmental factors to be considered when scheduling each task schedule are formulated as the state space, and the scheduling decision for each task is abstracted into the action space. A reward value is formulated with novelty, consisting of a utility function, a duration factor, and a penalty factor.
@@
    A DRL algorithm based on DQN is designed to learn the scheduling decision from the interaction with the environment. By conducting extensive simulations, the proposed task graph offloading algorithm is superior to existing strategies, in terms of reducing average makespan and deadline violation.
The remainder of this paper is organized as follows. Section 2 briefly surveys the related work. Section 3 presents the system model and Section 4 formulates the optimization problem of the task graph offloading in MEC. Section 5 gives the RL-based scheduling mechanism, and Section 6 describes the resource allocation based on RL. Then, the simulation evaluation and conclusions are provided in Section 7 and Section 8, respectively.
2 Related work
With advance of the Internet of Things (IoT) and Big data, smart mobile applications and IoT applications have proliferated rapidly IOT8930011 ; JIT2378 . Increasing researchers pay attention to the computing offloading in the edge computing. We briefly survey the related work in this section.
Task graph scheduling has always been an important issue in the heterogeneous computing systems li2013energy ; xu2014genetic . Some works extended the traditional task scheduling techniques in cloud computing to the edge computing. Zhang et al. zhang2018resource extended the partial critical path (PCP) to allocate computing resources for the task graph application. Our previous work liu2022auction regarded the computing resource allocation based on PCP for the task graph application as a trade. Then, we proposed an auction-based game to scheduling task graph applications when satisfying deadline constraints. Nevertheless, these works did not consider the MEC environment dynamic.
To adapt to the dynamic environment, many researchers have applied DRL to the computation offloading in dynamic MEC. He et al. he2017integrated improved the performance of computation offloading for vehicular networks by dynamic orchestration of network, caching, and computing resources. Ning et al. ning2019deep considered the variation of the channel state and the computation capability and proposed a DRL scheme to optimize traffic scheduling and resource allocation in vehicular networks. Chen et al. chen2018optimized took into account the time-varying network dynamics to investigate an optimal computation offloading policy in a sliced radio access network. They proposed a DRL-based offloading decision to maximize the long-term utility performance. However, these works did not take into account the dependency among tasks, that is, they focused on the coarse-grained task offloading.
There are also a few advances in utilizing DRL to further optimize the computing resource allocation of task graphs in the MEC system. Lu et al. lu2020optimization focused on the scenarios of the multi-service nodes and task graph applications in heterogeneous MEC. They proposed a fine-grained offloading scheme based on DRL to achieve the reduction of execution latency, monetary cost, energy consumption, as well as network usage. Song et al. song2022offloading considered the different dynamic preferences between MEC and users to investigate the offloading decision. They proposed a multi-objective optimization strategy based on DRL for task graph applications in MEC. Wang et al.wang2021dependent proposed a DRL-based task graph offloading scheme to make the task offloading plan, considering the variation of dependent constraints between tasks. These above-mentioned works did not take into account the variational computing capabilities of ECDs, though they improved the performance of task graphs.
Different from all the existing works above, this work investigates the fine-grained task graph offloading and proposes an offloading strategy based on DRL to adapt to the variational computing capabilities of ECDs.
@@ 3 System Mobel
A MEC system, as shown in Fig. 1, is comprised of many heterogeneous ECDs deployed at the edge of the mobile network. There exists a decision controller that resides in some ECDs in the system zhang2017resource . It can periodically collect the real-time status information broadcasted from all ECDs. In the decision controller, we design an intelligent agent of the reinforcement learning, unlike traditional task schedulers (e.g., YARN). The agent makes the decision intelligently for the task scheduling by observing changes in the MEC system.
@@
Let $\mathcal{M}$ express the set of ECDs in the MEC system and $M=|\mathcal{M}|$ denote the total number of ECDs. Then, we use $m$ ($\forall m\in\{1,\ldots,M\}$) to index the $m$-th ECD. These ECDs in $\mathcal{M}$ communicate with each other via wired links zhang2017resource . We indicate the transmission rate between any two ECDs as $B_{m,m^{\prime}}$, where $m\neq m^{\prime}$ and $\forall m,m^{\prime}\in\{1,\ldots,M\}$. Based on the literature zhang2015energy , $B_{m,m^{\prime}}$ can be estimated generally. Due to physical constraints, the computing resources of ECDs are limited. We assume that each ECD has only one processing element (PE) which provides the computing service for the task execution of the application following First Come First Service manner (FCFS). In practice, the processing capabilities of PEs are finite continuous values varying with time ning2019deep . Moreover, the processing capability state at the next moment is related to the previous moment. We assume that the processing capabilities of PEs are stabilized over a short period when they are within a level. Therefore, the processing capabilities of PEs can be discretized and quantized into several levels, similar to the literature ning2019deep . Then, let $\delta_{m}$ express the processing capability of ECD $m$ at a state, which can be measured in millions of instructions per second (MIPS). Each ECD can periodically broadcast the current status information, such as its own workload and the transmission rate with other ECDs. Since the work mainly focuses designing an intelligent algorithm to allocate the dependent tasks among different connected ECDs, we did not pay a special focus on the task scheduling over parallel processing elements. This simple model can also be easily extended to a more general case in which each ECD has multiple PEs with parallel computing capability by adding more virtual ECDs and neglecting the data transmission cost among them.
@@
In the MEC system shown in Fig. 1 each ECD can provide cellular communication services to many MUs in a specified area. We use $\mathcal{N}$ to indicate the set of all MUs that the MEC system can serve, and $N=|\mathcal{N}|$ denotes the total number of MUs. Thus, $n$ ($\forall n\in\{1,\ldots,N\}$) can index the $n$-th MU. Each MU needs to process an application composed of multiple dependent tasks. This proposal can also be easily extended to the scenarios in which a MU can process multiple applications by adding some new MUs that can only process one application to the system. As a result, $\mathcal{N}$ also represents the set of all applications and $n$ can index the $n$-th application [^1] In this paper, we will interchangeably use three terms, i.e., MU $n$, $n$-th MU and application $n$, to represent a user application. We can model the application as a direct acyclic graph (DAG) sundar2018offloading . Then, $\mathcal{G}_{n}(\mathcal{V}_{n},\mathcal{E}_{n})$ indicates a DAG corresponding to application $n$. $\mathcal{V}_{n}$ denotes the set of tasks. $\mathcal{E}_{n}$ is the set of data dependencies, called directed edges, among these tasks. For any task in $\mathcal{V}_{n}$, it is an atomic and indivisible component.
@@
MU $n$ expects to expedite the execution of its application before the deadline, so it needs to offload application data to the MEC system. To achieve this, MU $n$ must first send the offloading request to the ECD whose cellular communication service covers MU $n$. Let $m_{n}$ indicate the ECD that covers MU $n$, as shown in Fig. 1. $B_{n}^{m}$ denotes the transmission rate between ECD $m_{n}$ and MU $n$. After ECD $m_{n}$ accepts this request and receives the application data sent by MU $n$, the decision controller will extract the tasks ready for execution from ECD $m_{n}$ and push them into a ready queue $Q^{r}$. Whereafter, the decision controller can make the assignment for the tasks in $Q^{r}$ according to a scheduling strategy. Moreover, any task can only be executed on one ECD. The result data must send back to MU $n$ after all tasks of application $n$ are completed in the MEC system. As MU $n$ is located within the cellular signal coverage area of ECD $m_{n}$, the result data needs to be transmitted to MU $n$ via ECD $m_{n}$. For any MU $n$, it has the same network topology for task execution as all ECDs in this system can execute tasks offloaded from MU $n$. Hence, we only focus on the topology with one MU for simplicity, which is shown in Fig. 2. More MUs connected to the MEC system only increase the number of applications offloaded to this system, which does not affect this study. Built upon this topology with one MU, we can assume MU $n$ as the fictitious ECD which is the 00-th edge computing device in $\mathcal{M}$. To model the application data sent by MU $n$ and the result data received by MU $n$, we add two dummy tasks to each application in $\mathcal{N}$. Moreover, the two dummy tasks must be on this fictitious ECD corresponding to MU $n$. As a result, all tasks in application $n$ can be offloaded to the MEC system.
@@
Considering the above characteristics of the computing offloading, we use the triple elements $\{r_{n},d_{n},\mathcal{G}_{n}\}$ to indicate the application that MU $n$ offloads to the MEC system. Here, $r_{n}$ denotes the time of the offloading request sent by MU $n$ and $d_{n}$ is the deadline of application $n$. In the task graph $\mathcal{G}_{n}$, $I_{n}=|\mathcal{V}_{n}|$ expresses the number of tasks in application $n$. We indicate each task in $\mathcal{V}_{n}$ as $v_{ni}$ ($i\in\{0,\ldots,I_{n}\}$). Here, $i$ indexes the $i$-th task. For ease of expression, meanwhile, we use $v_{n0}$ and $v_{nI}$ to denote the two dummy tasks, respectively. For task $v_{ni}$, it has a workload $\rho_{ni}$, which is the number of instructions needed by the PE to process $v_{ni}$. Generally, $\rho_{ni}$ can be acquired by profile analytics venkataraman2016ernest . Accordingly, $\rho_{n0}$ and $\rho_{nI}$ are set to zero since $v_{n0}$ and $v_{nI}$ are the dummy tasks added to application $n$. Due to the dependency constraints among tasks, the output data from task $v_{ni}$ needs to be transferred to another task $v_{nj}$. Notice that, the dummy $v_{nI}$ has no the output data and the dummy $v_{n0}$ has no the input data. The data transfer between $v_{ni}$ and $v_{nj}$ forms a directed edge, denoted as $\varepsilon_{nij}\in\mathcal{E}_{n}$. If there exists the directed edge from $v_{ni}$ to $v_{nj}$, we call $v_{ni}$ the parent of $v_{nj}$. Accordingly, $v_{nj}$ is the child of $v_{ni}$. We use $p(v_{ni})$ and $c(v_{ni})$ to define the sets of $v_{ni}$’s parents and children, respectively. For the two dummy $v_{n0}$ and $v_{nI}$, especially, $p(v_{n0})$ and $c(v_{nI})$ are empty sets. Under the scheduling strategy made by the decision controller, the output data from a completed task need to be transferred to the ECD where the child tasks are located, due to the data dependencies between tasks. For example, task $v_{ni}$ is on ECD $2_{n}$ and $v_{ni}$’s child $v_{nj}$ is on ECD $3$ in Fig. 2. After ECD $2_{n}$ completes task $v_{ni}$, ECD $2_{n}$ needs to transmit the all output data to ECD $3$ via the communication link since $v_{ni}$’s child $v_{nj}$ is on ECD $3$. Let $e_{nij}$ denote the size of data transfer from $v_{ni}$ to $v_{nj}$. Especially, if both $v_{ni}$ and its child $v_{nj}$ on the same ECD $m$, the transmit data $e_{nij}$ from $v_{ni}$ can be delivered directly to $v_{nj}$ without going through the communication link. To help understand, the common symbols used throughout this paper are listed in Table 1.
Table 1: Frequently Used Symbols
Symbol
Definition
$M$, $n$
The total number of ECDs and the number of MUs
$B_{m,m^{\prime}}$
The transmission rate between $m$ and $m^{\prime}$
$m_{n}$
The ECD covering MU $n$
$B_{n}^{m}$
The transmission rate between ECD $m_{n}$ and MU $n$
$\hat{B}$
The maximum value of all $B_{m,m^{\prime}}$
$\delta_{m}$
The processing capability of ECD $m$
$\hat{\delta}$
The maximum value of all $\delta_{m}$
$\mathcal{V}_{n}$, $\mathcal{E}_{n}$
The sets of tasks and directed edges in DAG $\mathcal{G}_{n}$
$I_{n}$
The number of tasks in application $n$
$v_{ni}$
The $i$-th task in application $n$
$v_{n0}$ and $v_{nI}$
Two dummy tasks in application $n$
$\rho_{ni}$
A workload of task $v_{ni}$
$\varepsilon_{nij}$
The directed edge between $v_{ni}$ and $v_{nj}$
$p(v_{ni})$
The set of $v_{ni}$’s parents
$c(v_{ni})$
The set of $v_{ni}$’s children
$\textbf{x}_{ni}$
A strategy for task $v_{ni}$
$x_{ni}^{m}$
A binary variable denoting the scheduling plan
$E(\textbf{x}_{ni})$
The execution time of $v_{ni}$
$e_{nij}$
The transmit data between $v_{ni}$ to $v_{nj}$
$T(\textbf{x}_{ni},\textbf{x}_{nj})$
The data transfer time from $v_{ni}$ to $v_{nj}$ via the communication link
$F(\textbf{x}_{ni})$
The completion time of $v_{ni}$
$A(\textbf{x}_{nj})$
The arrival time of input data to $v_{ni}$
$F^{lct}_{ni}$
The latest completion time of $v_{ni}$
$\xi_{n}$
The list of tasks in application $n$ in ascending order
$\textbf{s}_{\tau}$
The state space in MDP at time step $\tau$
$\textbf{a}_{\tau}$
The action space in MDP at time step $\tau$
r The reward in MDP